OpenStudy (anonymous):

Evaluate the expression below when y = -3. Problem:2y + 7 over-4y - 10 Choices: A. -1 22 B. -13 2 C. 1 2 D. 1

OpenStudy (anonymous):

\[\frac{2y+7}{-4y-10}\] \[=\frac{2(-3)+7}{-4(-3)-10}\] \[=\frac{7-6}{12-10}\] \[=\frac{1}{2}\]

OpenStudy (anonymous):

(2(-3) + 7)/(-4(-3) -10) = (-6+7)/(12-10) = 1/2 --> the ans:)

OpenStudy (anonymous):

Thanks Guys !!